We departed Site 1223 at 0130 hr on 23 December 2001 and arrived in the vicinity of the H2O junction box (27°52.916'N, 141°59.504'W) at 0000 hr on 26 December to begin a seismic and 3.5-kHz echo sounder survey. All times are reported in ship local time, which is Universal Time Coordinated 9 hr at Site 1224. The 766-nmi voyage took 2.9 days at an average speed of 10.9 kt.
Following completion of the surveying at 0745 hr on 26 December, the JOIDES Resolution returned and positioned on proposed Site H2O-5 (Fig. F7) with Global Positioning System (GPS) navigation at 0845 hr on 26 December. Operations were suspended while waiting on weather (WOW) because of heave, pitch, roll, and wind up to 7.7 m, 5.2°, 4.5°, and 29 kt respectively. A total of 13.25 hr of WOW time occurred before drilling operations could proceed.
Prior to conducting drilling operations, the vibration isolated television (VIT) camera was launched to conduct a camera survey of the site for debris, while also conducting a survey with an echo sounder attached to the VIT frame to further delineate subsurface layers. The survey covered a 30 m x 30 m area, took 2.0 hr, and showed the site was flat, undisturbed, and free of debris and cables.
Drilling operations began when seafloor was tagged at 4966.1 meters below sea level (mbsl), or 4977 meters below rig floor (mbrf), at 1525 hr on 27 December. The jet-in test was performed to confirm a refusal depth for jetting-in the reentry cone with 20-in casing. At 12 to 13 mbsf a hard layer was encountered, although the test was suspect as the ship was experiencing 4 to 5 m heave at the time of the test. Following the test, we were again forced to WOW, this time for 13.75 hr.
Hole 1224A was spudded at 1455 hr on 28 December at 4977 mbrf. Core 200-1224A-1X was advanced 6 m downhole with no recovery; hence, we could not establish a precise mudline (Table T3). On Core 4X, drilling progress was slow when we got to hard rock, which at the time was thought to be chert or basaltic basement. Recovery of 1.24 m of red clay and pieces of basalt confirmed that we had penetrated basement near the bottom of the 5.5-m interval cored, or at ~28 mbsf. We attempted one more extended core barrel (XCB) core (5X), before switching to the motor-driven core barrel (MDCB) for one last short core. We pulled out of the hole and cleared the seafloor at 0530 hr on 29 December, ending Hole 1224A.
Overall we cored 32.2 m in Hole 1224A and recovered 1.67 m of core (5.19% recovery), with 32 m cored and 1.45 m recovered (4.53% recovery) with the XCB and 0.2 m cored and 0.22 m recovered (110% recovery) with the MDCB (Table T3).
Hole 1224B was spudded with the advanced piston corer (APC) at 0650 hr on 29 December, but only 0.2 m of core was recovered. The primary goal of APC coring was to establish the mudline; therefore, we offset to spud Hole 1224C. Hole 1224B officially ended at 0745 hr on 29 December after we pulled the bit up to clear the seafloor.
The bit was positioned at 4964.1 mbsl (4975 mbrf), and Hole 1224C was spudded with the APC at 0820 hr on 29 December. We recovered 6.53 m of core and established the mudline at 4967.1 mbsl (4978.0 mbrf). Having successfully determined the mudline, the bit was pulled clear of the seafloor at 0915 hr on 29 December, marking the end of Hole 1224C.
Operations were delayed by bad weather, which included maximum heave, pitch, and roll of 6.3 m, 2.4°, and 8.1°, respectively, with winds up to 44 kt. Total time WOW was 16.0 hr, with operations beginning again at 0115 hr on 30 December.
A second jet-in test was deemed necessary to confirm the depth of penetration for the 20-in surface casing, which would be run with the reentry cone. A wash barrel was dropped and the bottom-hole assembly (BHA) was jetted-in to 4996.1 mbsl (5007 mbrf), ~29 mbsf, with no obstructions encountered, unlike the first jet-in test. The drill string was pulled out of the hole, with the bit clearing the rotary table at 1430 hr on 30 December.
The reentry cone was positioned over the moonpool doors, and the casing string was partially assembled at 1830 hr on 30 December 2001. Poor weather conditions and the associated large heave, roll, and pitch forced us to delay operations until 1715 hr on 1 January 2002, a loss of 46.75 hr.
With weather conditions improving, the reentry cone and ~25 m of 20-in casing were assembled and lowered through the moonpool at 2335 hr on 1 January. Hole 1224D was spudded at 1220 hr on 2 January. It took only 24 min to jet the 20-in casing string down to 5003.47 mbrf (25.47 mbsf) and set the reentry cone. VIT observation of the reentry cone confirmed that it was in a satisfactory position. The bit cleared the seafloor at 1315 hr on 2 January and the pipe was tripped back to the rig floor, with the jet-in BHA and bit clearing the rotary table at 0200 hr on 3 January.
Hole 1224D was reentered with an RCB bit at 1837 hr on 3 January, with coring beginning at 25.5 mbsf. Coring progressed down to 59 mbsf, with several delays caused by the poor weather conditions, adding up to another 26.0 hr of WOW. The marine forecast was for continued poor weather for our operating area, with very strong low-pressure systems to the west and north and large swells. It was therefore decided to prepare to take advantage of any weather window by tripping the drill string and changing to the 14.75-in bit and BHA. This would allow us to open the cored hole when a more appropriate weather window was available and be in position to run 10.75-in casing. At 0045 hr on 7 January we started to trip the pipe, with the bit clearing the rotary at 1100 hr on 7 January.
Overall we cored 33.5 m in Hole 1224D and recovered 15.65 m of core (46.72% recovery) with the RCB coring system (Table T3).
After tripping the pipe with the 14.75-in bit to 4790 mbrf at 2345 hr on 7 January, operations were again put on hold while WOW for 19.0 hr. Operations resumed at 1845 hr on 8 January after Hole 1224D was reentered.
We reamed the hole to 64.7 mbsf before drilling difficulties halted penetration. This was sufficiently close to the planned drilling depth of 67 mbsf, so we ceased drilling at 0730 hr on 11 January, for a total depth in Hole 1224D of 64.7 mbsf. When the drill string was pulled to the surface, to the surprise of all, the bit had been left in the hole, thus explaining the drilling difficulties. The bit appeared to have been sheared off.
Starting at 2030 hr on 11 January, the drill crew began assembling the 10.75-in casing string, which consisted of five joints of 10.75-in (40.5 lb/ft) casing. Hole 1224D was reentered by the casing string at 1336 hr on 12 January. We noted during reentry that the reentry cone and skirt had settled by ~1.7 m below the original mudline. The casing string was run down and landed with the base at 5036.47 mbrf (58.47 mbsf) on 1515 hr on 12 January. The casing was cemented with 18.8 bbl of 15.5 ppg Class G cement. The first attempt to release from the casing hanger failed and resulted in the 10.75-in casing hanger being pulled up above the reentry cone. The casing hanger was landed again in the 20-in casing hanger and this time the 10.75-in hanger released at 1715 hr on 12 January. The pipe was tripped up, with the running tool clearing the rotary table at 0530 hr on 13 January.
The BHA was assembled with a RCB bit and run down to 4388.89 mbrf in preparation for coring in Hole 1224E. Before starting Hole 1224E, we reentered Hole 1224D to ensure that the casing and cement were properly installed.
The JOIDES Resolution was offset 15 m to the southwest, and Hole 1224E was spudded at 1840 hr on 13 January at 4978 mbrf. We washed down the first 8 m and then took two punch, or push, cores (1R and 2R), which were acquired by lowering the RCB bit through the soft sediments without rotating the bit. Both cores sustained substantial drilling disturbance, but we were able to recover 10.52 m of sediment core in a 19.2-m-long interval from 8.0 to 27.1 mbsf, where recovery was virtually absent in the other holes.
Coring penetrated from 27.1 to 36.7 mbsf for Core 3R. Basement was tagged at 27.7 mbsf during coring. Recovery consisted of basaltic basement underlying a 5-cm-thick piece of hyaloclastite, into which basalt glass and clay pieces had been incorporated. This likely is the top few centimeters of the basement. After completing coring on Core 3R, the bit was pulled up by one stand of drill pipe to connect another joint of pipe. This placed the bit above the sediment/basement contact. After making the connection, the driller was unable to reenter the basement hole. After 1 hr of attempting to find the hole by rotating the bit on bottom, a new hole (1224F) was started.
Overall we cored 28.7 m in Hole 1224E and recovered 14.91 m of core (51.95% recovery) with the RCB coring system (Table T3).
The start of Hole 1224F is somewhat of an anomaly in the ODP nomenclature, since the bit never pulled totally out of Hole 1224E, but it did pull out of the basement portion of Hole 1224E. The distance between Holes 1224E and 1224F is likely no more than about a meter. In any case, we began penetrating basement again at 1630 hr on 14 January in Hole 1224F.
For all the bad weather we had previously, we were due a good spell. Thus, coring proceeded without interruption except for the occasional wiper trip and one trip to replace the knobby joints with drill pipe. During the latter trip, which started at 2315 hr on 17 January after recovery of Core 11R, the bit was inadvertently pulled above the basement/sediment contact. The driller worked the drill string up and down with rotation in an attempt to reenter Hole 1224F. Instead, Hole 1224E was reentered five times before finally the bit went back into Hole 1224F. RCB coring proceeded after washing ~11 m of soft fill from the bottom of hole. Cores continued to be cut at a rate of ~68 hr/core, which was roughly twice as fast as cores cut from near the top of the basement. No core was recovered in Core 16R. The bit deplugger was run to remove potential obstructions, but Core 17R also had no recovery. Owing to time limitation, coring in Hole 1224F ended and preparations for logging began.
Overall in Hole 1224F, we penetrated 174.5 m, cored 146.8 m, and recovered 37.7 m of core (25.68% recovery) with the RCB coring system (Table T3).
The bit was released in the bottom of the hole at 2320 hr on 19 January. The hole was then displaced with 75 bbl of sepiolite mud. A free-fall funnel (FFF) was launched at 0442 hr on 20 January to facilitate reentries into Hole 1224F on future scientific experiments.
At 0730 hr on 20 January, the triple combo tool was prepared to run downhole. The tool reached 5152 mbrf, which is only 0.5 m off the bottom of the hole. The first logging run was completed, and the tool was through the rotary table at 1520 hr on 20 January. For each logging run, the base of the pipe was lowered to 49.9 mbsf initially. As each run was made uphole, the pipe was pulled up from 49.9 mbsf to 34.5 mbsf to increase the open-hole interval for logging.
The second logging run was with the Formation MicroScanner/dipole sonic imager (FMS/DSI) tool. Three passes of this string were run uphole at 275 m/hr from the bottom of the hole to the basement contact (27.7 mbsf). The second logging run was completed, and the tool cleared the rotary table at 0525 on 21 January.
We had planned to test the three-component well seismic tool (WST-3) if time and weather conditions permitted. While testing the tool and air gun, three problems were found: a faulty circuit in the blast hydrophone of the air gun, an air leak from the air gun, and the WST-3 telemetry worked intermittently. The experiment was thus terminated because there was insufficient time to attempt to fix the problems and complete the planned shooting program to the WST-3. The time constraint on the logging program was determined by the departure time required to make the San Diego port call. The WST was back through the rotary table at 0945 hr on 21 January.
The VIT was launched starting at 1030 hr 21 January to observe the FFF at the top of Hole 1224F. A large hole was observed in the seafloor from circulating the cuttings out of the hole. As a result the top of the FFF was observed at ~4980.5 mbrf (2.5 mbsf) with the three buoys just below the mudline, secured to the FFF by a 5/32 -in steel cable. The end of the casing on the FFF is estimated to be at 6.2 mbsf.
The open end of the drill pipe cleared the seafloor and FFF at 1238 hr on 21 January. The VIT was recovered at 1445 hr, and the BHA cleared the rotary table at 2355 hr on 21 January, completing activity at Site 1224.
Next Section | Table of Contents